# TCJA-SNN: Temporal-Channel Joint Attention for Spiking Neural Networks

### v2

@@ Abstract
Spiking Neural Networks (SNNs) are attracting widespread interest due to their biological plausibility, energy efficiency, and powerful spatio-temporal information representation ability. Given the critical role of attention mechanisms in enhancing neural network performance, the integration of SNNs and attention mechanisms exhibits tremendous potential to deliver energy-efficient and high-performance computing paradigms. In this paper, we present a novel Temporal-Channel Joint Attention mechanism for SNNs, referred to as TCJA-SNN. The proposed TCJA-SNN framework can effectively assess the significance of spike sequence from both spatial and temporal dimensions. More specifically, our essential technical contribution lies on: 1) We employ the squeeze operation to compress the spike stream into an average matrix. Then, we leverage two local attention mechanisms based on efficient 1-D convolutions to facilitate comprehensive feature extraction at the temporal and channel levels independently. 2) We introduce the Cross Convolutional Fusion (CCF) layer as a novel approach to model the inter-dependencies between the temporal and channel scopes. This layer effectively breaks the independence of these two dimensions and enables the interaction between features. Experimental results demonstrate that the proposed TCJA-SNN outperforms the state-of-the-art (SOTA) by up to 15.7% classification accuracy on all standard static and neuromorphic datasets, including Fashion-MNIST, CIFAR10-DVS, N-Caltech 101, and DVS128 Gesture. Furthermore, we effectively apply the TCJA-SNN framework to image generation tasks by leveraging a variation autoencoder. To the best of our knowledge, this study is the first instance where the SNN-attention mechanism has been employed for high-level classification and low-level generation tasks. Notably, our approach has achieved state-of-the-art performance in both domains, establishing a significant advancement in the field. Our implementation codes are available at https://github.com/ridgerchu/TCJA.
@@ I Introduction
Spiking Neural Networks (SNNs) have emerged as a promising research area, offering lower energy consumption and superior robustness compared to conventional Artificial Neural Networks (ANNs) [1, 2]. These characteristics make SNNs highly promising for temporal data processing and power-critical applications [1, 3]. In recent years, significant progress has been made by incorporating backpropagation into SNNs [4, 5, 6, 7, 8, 3], which allows the integration of various ANN modules into SNN architectures, including batch normalization blocks [9] and residual blocks [10]. By leveraging these ANN-based methods, it becomes possible to train large-scale SNNs while preserving the energy efficiency associated with SNN’s binary spiking nature.
Despite significant progress, SNNs have yet to fully exploit the superior representational capability of deep learning, primarily due to their unique training mode, which struggles to model complex channel-temporal relationships effectively. Addressing this limitation, Zheng et al. [9] introduced a batch normalization method for the temporal dimension, overcoming issues of gradient vanishing and threshold-input balance. On the other hand, Wu et al. [11] proposed a method named NeuNorm to address the channel-wise challenges. NeuNorm includes an auxiliary neuron that adjusts the stimulus strength generated by the preceding layer, enhancing performance while mimicking the activity of the retina and nearby cells for added bio-plausibility. However, existing methods handle temporal and channel information separately, leading to limited joint information extraction. Given that SNNs reuse network parameters at each time step, there exists untapped potential for recalibration at both the temporal and channel dimensions. Especially, TA-SNN proposed by Yao et al. [12].
Previous studies in ANNs [13, 14] have often utilized the attention mechanism as a means to address the challenges posed by multidimensional dynamic problems. The attention mechanism, inspired by human cognitive processes, enables the selective focus on relevant information while disregarding irrelevant data. This approach has shown promise in the realm of SNNs and merits further exploration [12]. For instance, in the domain of neuroscience, an attention-based Spike-Timing-Dependent Plasticity (STDP) SNN was proposed by Bernert et al.[15] to tackle the spike-sorting problem. Furthermore, Yao et al.[12] incorporated a channel attention block into the temporal-wise input of an SNN, as depicted in Fig. 1-a, enabling the assessment of frame significance during training and the exclusion of irrelevant frames during inference. Despite employing attention solely in the temporal dimension, this attention mechanism significantly improves the network’s performance.
@@
    Experimental results show that TCJA-SNN outperforms previous methods on static and neuromorphic datasets for classification tasks. It also performs well in generation tasks. Notably, our model achieves an impressive 82.5% test accuracy on the N-Caltech 101 dataset, surpassing the previous state-of-the-art accuracy of 66.8% by a substantial margin.
@@ II-A Training Techniques for Spiking Neural Networks
In recent years, the direct application of various ANNs algorithms for training deep SNNs, including gradient-descent-based methods, has gained traction. However, the non-differentiability of spikes poses a significant challenge. The Heaviside function, commonly used to trigger spikes, has a derivative that is zero everywhere except at the origin, rendering gradient-based learning infeasible. To overcome this obstacle, a commonly employed solution is the surrogate gradient descent method [16, 17, 18, 19, 20, 9, 21].
@@ II-B Attention Mechanism in Convolutional Neural Networks
In the realm of ANNs, the Squeeze and Excitation (SE) block, introduced by Hu et al. [13], has proven to be a highly effective module for enhancing representation. The SE block can be seamlessly incorporated into a network, requiring only a minimal increase in parameters to recalibrate channel information. By employing squeezing and fully connecting operations, it allows the network to learn a trainable scale factor for each channel. This recalibration process significantly improves the discriminative power of individual channels. Recently, Yao et al. [12] extended the application of the SE block to SNNs by formulating a temporal-wise attention mechanism. This innovative approach enables SNNs to identify critical temporal frames of interest without being adversely affected by noise or interference. By incorporating temporal-wise attention, the proposed technique achieves state-of-the-art performance across various datasets. This accomplishment serves as compelling evidence for the immense potential of attention mechanisms within SNNs. The utilization of SE blocks and the introduction of temporal-wise attention represent significant advancements in the field of SNNs. These techniques not only enhance the representation capability of SNNs but also offer insights into effectively leveraging attention mechanisms for improved performance. In the following sections, we aim to explore and further leverage these attention mechanisms to improve the performance of SNNs and unlock their full potential in complex temporal data processing tasks.
@@ II-C Motivation
Based on the aforementioned analysis, the utilization of a temporal-wise attention mechanism in SNNs has exhibited substantial progress in effectively processing time-related data streams. Moreover, it has been observed in both biological neural networks [25] and ANNs [13] that recalibrating channel features within convolutional layers hold considerable potential for enhancing performance. However, the existing SNNs-based works only process the data with either temporal or channel dimensions, thereby constraining the capacity for joint feature extraction. To present the correlation between time steps and channels, we visualize the input frame and several proximity channels output from the first 2-D convolutional layer shown in Fig. 2. As the circles indicate, a similar firing pattern can be distinguished from the surrounding time steps and channels. To fully use this associated information, we propose the TCJA module, a novel approach for modeling temporal and channel-wise frame correlations. Furthermore, considering the inevitable increases in the model parameters caused by the attention mechanism, we attempt to adopt the 1-D convolution operation to gain a reasonable trade-off between model performance and parameters. Furthermore, existing SNN attention mechanisms primarily prioritize classification tasks, neglecting the needs of generation tasks. Our goal is to introduce an attention mechanism that can proficiently handle both classification and generation tasks, thereby establishing a universal attention mechanism for SNNs.
Figure 2: Correlation between proximity time steps and channels. The top row is the input frame. Each figure in the nine-pattern grid of the bottom row denotes a channel output from the first 2-D convolutional layer. It is clear that a significant correlation exists in channels with varying time steps, motivating us to merge the temporal and channel information.
@@ III-A Leaky Integrate and Fire Model
As a mainstream neuron model, LIF-based SNN models can be trained directly using surrogate gradient methods [6, 20] to attain state-of-the-art (SOTA) performance [29, 23, 12]. Moreover, the LIF model is well-suited to common machine-learning frameworks because it allows forward and backward propagation along spatial and temporal dimensions. In our method, the parameters of the LIF model are set as follows:$\tau=2$, $\boldsymbol{V}_{reset}=0$, and $\boldsymbol{V}_{threshold}=1$.
@@ III-B Temporal-Channel Joint Attention (TCJA)
Figure 3: The growth curve of parameters between Fully-Connected (FC) layer and Temporal-Channel Joint Attention (TCJA) layer when $C=64$.
As mentioned above, we contend that the frame at the current time step exhibits a significant correlation with its neighboring frames in both the channel and temporal dimensions. This correlation opens up the possibility of employing a mechanism to establish a connection between these two dimensions. Initially, we employed a fully-connected layer to establish the correlation between the temporal and channel information, as it provides the most direct and prominent connection between these dimensions. However, as the number of channels and time steps increases, the number of parameters grows rapidly with a ratio of $T^{2}\times C^{2}$, as illustrated in Fig.  3. our subsequent attempt involved utilizing a 2-D convolutional layer for building this attention mechanism. Nonetheless, this solution presented its own constraint, as the fixed kernel size limited the receptive field. For this reason, it is necessary to decrease the number of parameters while increasing the receptive field. In Sec. IV-G5, we provide a detailed theoretical analysis of the receptive field.
@@
Figure 4: The Framework of SNN with TCJA module. The information flown in SNNs is a form of spike sequence along both temporal and spatial dimensions. In temporal-wise, the spiking neuron feed-forward in membrane potential ($\boldsymbol{V}$) and spike ($\boldsymbol{S}$) as the Eq. 2, and backpropagation with the surrogate function. In spatial-wise, data flows between layers as ANN. TCJA module first compresses the information both temporal-wise and spatial-wise, then apply TLA and CLA to establish the relationship in both temporal and channel dimensions and blend them by CCF layer.
@@ III-B3 Channel-Wise Local Attention (CLA)
where $K_{C}$ ($K_{C}$ <$C$) represents the size of the convolution kernel, and $E_{(n,i)}^{m}$ is a learnable parameter, representing the $m$-th parameter of the $i$-th channel when performing $T$-channel 1-D convolution on $n$-th column of $\mathcal{Z}$. $\mathcal{C}\in\mathbb{R}^{C\times T}$ is the attention score matrix after CLA mechanism. To ensure that the input and output dimensions are consistent, we employ the same padding technique on TLA and CLA.
@@ III-B4 Cross Convolutional Fusion (CCF)
After TLA and CLA operations, we get the temporal (TLA matrix $\mathcal{T}$) and channel (CLA matrix $\mathcal{C}$) saliency scores of the input frame and its adjacent frames, respectively. Next, to learn the correlation between temporal and channel frames in tandem, we propose a cross-domain information fusion mechanism, i.e., the CCF layer. The goal of CCF is to calculate a fusion information matrix $\mathcal{F}$, and $\mathcal{F(i,j)}$ is used to measure the potential correlation between the $i$-th channel of the $j$-th input temporal frame and other frames. Specifically, the joint relationship between frames can be obtained by performing an element-wise multiplication of $\mathcal{T}$ and $\mathcal{C}$ as follows:
@@ III-C Training Framework
We integrate the TCJA module into the existing benchmark SNNs and propose the TCJA-SNN. Since the process of neuron firing is non-differentiable, we utilize the derived ATan surrogate function $\sigma^{\prime}(x)=\frac{\alpha}{2(1+(\frac{\pi}{2}\alpha x)^{2})}$ and the derived triangle-like surrogate function $\epsilon^{\prime}(x)=\frac{1}{\gamma^{2}}\text{max}(0,\gamma-|{x}-1|)$ for backpropagation, which is proposed by [23] and [24], respectively. In our method, the Spike Mean-Square-Error (SMSE) [16, 23] is chosen as the loss function, which can be expressed as:
@@
where $T$ is the total simulation time, $\mathcal{L}_{\text{CE}}$ denotes the cross-entropy loss, $s$ is the network output, and $g$ represents the target label. To estimate the classification accuracy, we define the predicted label $l_{p}$ as the index of the neuron with the highest firing rate $l_{p}=\mathop{\max_{i}}\frac{1}{T}\sum_{t=0}^{T-1}s_{t,i}$. Since the TCJA module simply utilizes the 1-D convolutional layer and Sigmoid function, it can be effortlessly introduced into the current network architecture as a plug-and-play module without adjusting to backpropagation.
@@ IV Experiments
We evaluate the classification performance of TCJA-SNN on both neuromorphic datasets (CIFAR10-DVS, N-Caltech 101, and DVS128 Gesture) and static datasets (Fashion-MNIST). Note that all neuromorphic datasets are collected from the event sensor. To verify the effectiveness of the proposed method, we integrate the TCJA module into several architectures [29, 23] with competitive performance to see if the integrated architecture can generate significant improvement.

### v3

@@ Abstract
Spiking Neural Networks (SNNs) are attracting widespread interest due to their biological plausibility, energy efficiency, and powerful spatio-temporal information representation ability. Given the critical role of attention mechanisms in enhancing neural network performance, the integration of SNNs and attention mechanisms exhibits tremendous potential to deliver energy-efficient and high-performance computing paradigms. In this paper, we present a novel Temporal-Channel Joint Attention mechanism for SNNs, referred to as TCJA-SNN. The proposed TCJA-SNN framework can effectively assess the significance of spike sequence from both spatial and temporal dimensions. More specifically, our essential technical contribution lies on: 1) We employ the squeeze operation to compress the spike stream into an average matrix. Then, we leverage two local attention mechanisms based on efficient 1-D convolutions to facilitate comprehensive feature extraction at the temporal and channel levels independently. 2) We introduce the Cross Convolutional Fusion (CCF) layer as a novel approach to model the inter-dependencies between the temporal and channel scopes. This layer effectively breaks the independence of these two dimensions and enables the interaction between features. Experimental results demonstrate that the proposed TCJA-SNN outperforms the state-of-the-art on all standard static and neuromorphic datasets, including Fashion-MNIST, CIFAR10, CIFAR100, CIFAR10-DVS, N-Caltech 101, and DVS128 Gesture. Furthermore, we effectively apply the TCJA-SNN framework to image generation tasks by leveraging a variation autoencoder. To the best of our knowledge, this study is the first instance where the SNN-attention mechanism has been employed for high-level classification and low-level generation tasks. Our implementation codes are available at https://github.com/ridgerchu/TCJA.
@@ I Introduction
Spiking Neural Networks (SNNs) have emerged as a promising research area, offering lower energy consumption and superior robustness compared to conventional Artificial Neural Networks (ANNs) [1, 2]. These characteristics make SNNs highly promising for temporal data processing and power-critical applications [1, 3]. In recent years, significant progress has been made by incorporating backpropagation into SNNs [4, 5, 6, 7, 8, 3, 9, 10], which allows the integration of various ANN modules into SNN architectures, including batch normalization blocks [11] and residual blocks [12]. By leveraging these ANN-based methods, it becomes possible to train large-scale SNNs while preserving the energy efficiency associated with SNN’s binary spiking nature.
Despite significant progress, SNNs have yet to fully exploit the superior representational capability of deep learning, primarily due to their unique training mode, which struggles to model complex channel-temporal relationships effectively. To address this limitation, Zheng et al. [11] introduced a batch normalization method for the temporal dimension, overcoming issues of gradient vanishing and threshold-input balance. On the other hand, Wu et al. [13] proposed a method named NeuNorm to address the channel-wise challenges. NeuNorm includes an auxiliary neuron that adjusts the stimulus strength generated by the preceding layer, enhancing performance while mimicking the activity of the retina and nearby cells for added bio-plausibility. However, existing methods handle temporal and channel information separately, leading to limited joint information extraction. Given that SNNs reuse network parameters at each time step, there exists untapped potential for recalibration at both the temporal and channel dimensions. Especially, TA-SNN proposed by Yao et al. [14].
Previous studies in ANNs [15, 16] have often utilized the attention mechanism as a means to address the challenges posed by multidimensional dynamic problems. The attention mechanism, inspired by human cognitive processes, enables the selective focus on relevant information while disregarding irrelevant data. This approach has shown promise in the realm of SNNs and merits further exploration [14]. For instance, in the domain of neuroscience, an attention-based Spike-Timing-Dependent Plasticity (STDP) SNN was proposed by Bernert et al.[17] to solve the spike-sorting problem. Furthermore, Yao et al.[14] incorporated a channel attention block into the temporal-wise input of an SNN, as depicted in Fig. 1-a, enabling the assessment of frame significance during training and the exclusion of irrelevant frames during inference. Despite employing attention solely in the temporal dimension, this attention mechanism significantly improves the network’s performance.
@@
    Experimental results show that TCJA-SNN outperforms previous methods on static and neuromorphic datasets for classification tasks. It also performs well in generation tasks.
@@ II-A Training Techniques for Spiking Neural Networks
In recent years, the direct application of various ANNs algorithms for training deep SNNs, including gradient-descent-based methods, has gained traction. However, the non-differentiability of spikes poses a significant challenge. The Heaviside function, commonly used to trigger spikes, has a derivative that is zero everywhere except at the origin, rendering gradient-based learning infeasible. To overcome this obstacle, the commonly employed solutions are ANN-to-SNN [18, 19, 20] and the surrogate gradient descent method [21, 22, 23, 24, 25, 26, 27, 28].
@@ II-B Attention Mechanism in Convolutional Neural Networks
In the realm of ANNs, the Squeeze and Excitation (SE) block, introduced by Hu et al. [15], has proven to be a highly effective module for enhancing representation. The SE block can be seamlessly incorporated into a network, requiring only a minimal increase in parameters to recalibrate channel information. By employing squeezing and fully connecting operations, it allows the network to learn a trainable scale factor for each channel. This recalibration process significantly improves the discriminative power of individual channels. Recently, Yao et al. [14] extended the application of the SE block to SNNs by formulating a temporal-wise attention mechanism. This innovative approach enables SNNs to identify critical temporal frames of interest without being adversely affected by noise or interference. By incorporating temporal-wise attention, the proposed technique achieves state-of-the-art performance across various datasets. This accomplishment serves as compelling evidence for the immense potential of attention mechanisms within SNNs. The utilization of SE blocks and the introduction of temporal-wise attention represent significant advancements in the field of SNNs. These techniques not only enhance the representation capability of SNNs but also offer insights into effectively leveraging attention mechanisms for improved performance. In the Sec. III, we aim to explore and further leverage these attention mechanisms to improve the performance of SNNs and unlock their full potential in complex temporal data processing tasks.
@@ II-C Motivation
Based on the aforementioned analysis, the utilization of a temporal-wise attention mechanism in SNNs has exhibited substantial progress in effectively processing time-related data streams. Moreover, it has been observed in both biological neural networks [32] and ANNs [15] that recalibrating channel features within convolutional layers hold considerable potential for enhancing performance. Nevertheless, the existing SNNs-based works only process the data with either temporal or channel dimensions, thereby constraining the capacity for joint feature extraction. To illustrate the relationship between temporal steps and channel dimensions, we provide a visual representation. This is achieved by displaying the input frame alongside several adjacent channel outputs, which originate from the initial 2-D convolutional layer, as demonstrated in Fig. 2. As the circles indicate, a similar firing pattern can be distinguished from the surrounding time steps and channels. To fully use this associated information, we propose the TCJA module, a novel approach for modeling temporal and channel-wise frame correlations. Furthermore, considering the inevitable increases in the model parameters caused by the attention mechanism, we attempt to adopt the 1-D convolution operation to gain a reasonable trade-off between model performance and parameters. Furthermore, existing SNN attention mechanisms primarily prioritize classification tasks, neglecting the needs of generation tasks. Our goal is to introduce an attention mechanism that can proficiently handle both classification and generation tasks, thereby establishing a universal attention mechanism for SNNs.
Figure 2: Correlation between proximity time steps and channels. The top row is the input frame selected from DVS128 Gesture dataset. Each figure in the nine-pattern grid of the bottom row denotes a channel output from the first 2-D convolutional layer. It is clear that a significant correlation exists in channels with varying time steps, motivating us to merge the temporal and channel information.
@@ III-A Leaky Integrate and Fire Model
As a mainstream neuron model, LIF-based SNN models can be trained directly using surrogate gradient methods [24] to attain state-of-the-art (SOTA) performance [36, 30, 14]. Moreover, the LIF model is well-suited to common machine-learning frameworks because it allows forward and backward propagation along spatial and temporal dimensions. In our method, the parameters of the LIF model are set as follows:$\tau=2$, $\boldsymbol{V}_{reset}=0$, and $\boldsymbol{V}_{threshold}=1$.
@@ III-B Temporal-Channel Joint Attention (TCJA)
Figure 3: The growth curve of parameters between Fully-Connected (FC) layer and TCJA layer when channel size $C=64$.
As mentioned above, we contend that the frame at the current time step exhibits a significant correlation with its neighboring frames in both the channel and temporal dimensions. This correlation opens up the possibility of employing a mechanism to establish a connection between these two dimensions. Initially, we employed a fully-connected layer to establish the correlation between the temporal and channel information, as it provides the most direct and prominent connection between these dimensions. However, as the number of channels and time steps increases, the number of parameters grows rapidly with a ratio of $T^{2}\times C^{2}$, as illustrated in Fig.  3. our subsequent attempt involved utilizing a 2-D convolutional layer for building this attention mechanism. Nevertheless, this approach encountered a limitation due to the fixed kernel size, which restricts the receptive field to a confined local area. In conventional CNNs, augmenting the number of layers can expand the receptive field [37, 38]. However, within the context of attention mechanisms, the feasibility of layer stacking, analogous to convolutional networks, is constrained, thereby limiting the receptive field when employing 2D convolutions. For this reason, it is necessary to decrease the number of parameters while increasing the receptive field. In Sec. IV-G5, we provide a detailed theoretical analysis of the receptive field.
@@
Figure 4: The Framework of SNN with TCJA module. In SNNs, information is transmitted in the form of spike sequences, encompassing both temporal and spatial dimensions. In temporal-wise, the spiking neuron with a threshold feed-forward in membrane potential ($\boldsymbol{V}$) and spike ($\boldsymbol{S}$) as the Eq. 2, and backpropagation with the surrogate function. In spatial-wise, data flows between layers as ANN. The TCJA module operates by initially compressing information along both temporal and spatial dimensions, then apply TLA and CLA to establish the relationship in both temporal and channel dimensions and blend them by CCF layer.
@@ III-B3 Channel-Wise Local Attention (CLA)
where $K_{C}$ ($K_{C}$ <$C$) represents the size of the convolution kernel, and $E_{(n,i)}^{m}$ is a learnable parameter, representing the $m$-th parameter of the $i$-th channel when performing $T$-channel 1-D convolution on $n$-th column of $\mathcal{Z}$. $\mathcal{C}\in\mathbb{R}^{C\times T}$ is the attention score matrix after CLA mechanism.
To maintain dimensional consistency between the input and output, a “same padding” technique is employed in both the TLA and CLA mechanisms. This padding strategy ensures that the output dimension matches the input dimension by adding an appropriate number of zeros to the input data. Specifically, this technique involves padding the input array with zeros on both sides, where the number of zeros added is determined based on the kernel size and the stride.
@@ III-B4 Cross Convolutional Fusion (CCF)
After TLA and CLA operations, we get the temporal (TLA matrix $\mathcal{T}$) and channel (CLA matrix $\mathcal{C}$) saliency scores of the input frame and its adjacent frames, respectively. Next, to learn the correlation between temporal and channel frames in tandem, we propose a cross-domain information fusion mechanism, i.e., the CCF layer. The goal of CCF is to calculate a fusion information matrix $\mathcal{F}$, and $\mathcal{F(\textit{i},\textit{j})}$ is used to measure the potential correlation between the $i$-th channel of the $j$-th input temporal frame and other frames. Specifically, the joint relationship between frames can be obtained by performing an element-wise multiplication of $\mathcal{T}$ and $\mathcal{C}$ as follows:
@@ III-C Training Framework
We integrate the TCJA module into the existing benchmark SNNs and propose the TCJA-SNN. Since the process of neuron firing is non-differentiable, we utilize the derived ATan surrogate function $\sigma^{\prime}(x)=\frac{\alpha}{2(1+(\frac{\pi}{2}\alpha x)^{2})}$ and the derived triangle-like surrogate function $\epsilon^{\prime}(x)=\frac{1}{\gamma^{2}}\text{max}(0,\gamma-|{x}-1|)$ for backpropagation, which is proposed by [30] and [31], respectively. This latter function is particularly applied in the TCJA-TET-SNN, in alignment with the default surrogate function specification for TET-based architectures. In our method, the Spike Mean-Square-Error (SMSE) [21, 30] is chosen as the loss function, which can be expressed as:
@@
where $T$ is the total simulation time, $\mathcal{L}_{\text{CE}}$ denotes the cross-entropy loss, $s$ is the network output, and $g$ represents the target label. The cross entropy loss here can be represented by:
$\mathcal{L}_{\text{CE}}(p,y)=-\sum_{c=1}^{M}y_{o,c}\log(p_{o,c})$ (9)
where $M$ is the number of classes, $y_{o,c}$ is a binary indicator (0 or 1) if class label $c$ is the correct classification for observation $o$, $p_{o,c}$ is the predicted probability of observation $o$ being of class $c$.
To estimate the classification accuracy, we define the predicted label $l_{p}$ as the index of the neuron with the highest firing rate $l_{p}=\mathop{\max_{i}}\frac{1}{T}\sum_{t=0}^{T-1}s_{t,i}$. Since the TCJA module simply utilizes the 1-D convolutional layer and Sigmoid function, it can be effortlessly introduced into the current network architecture as a plug-and-play module without adjusting to backpropagation.
@@ IV Experiments
We evaluate the classification performance of TCJA-SNN on both neuromorphic datasets (CIFAR10-DVS, N-Caltech 101, and DVS128 Gesture) and static datasets (Fashion-MNIST, CIFAR10, CIFAR100). Note that all neuromorphic datasets are collected from the event sensor. To verify the effectiveness of the proposed method, we integrate the TCJA module into several architectures [36, 30] with competitive performance to see if the integrated architecture can generate significant improvement.